The perimeter of a rectangle is 50 in. and the area is 144 in^2. Find the dimensions of the rectangle.
Answer by MithrilKnights(8)

We want to convert to equations
The perimeter = 2l + 2w = 50
So, l + w = 25
The area = l * w = 144
Finding w in terms of l we get w = 25-l
Plugging this into the second equation, we get (25-l)(l) = 144
We get -l^2 +25l -144 = 0
Multiplying by -1 we get l^2 -25l +144
to solve this, we would find the two numbers adding to -25 and product of 144.
The two numbers are -16, and - 9
So, l is either 9 or 16. And w is either 9 or 16.
So, l is 16 since it is bigger, and w is 9
The dimensions of the rectangle are 16 by 9